Another student has heard how smart you are and needs help finding the measure of 2 angles after being told only that Sine = Cos and that angle has (4x + 10) degrees and α β α that angle has (10x + 10) degrees. Please help him to find the measure of each angle and β add an explanation of the rules applied here and the steps taken so that this student can do this on their own next time.
1
Expert's answer
2020-03-21T14:09:13-0400

It is stated in the task sin=cos

We do not know if it means sin"\alpha" = cos "\alpha" , sin "\beta" = cos "\beta" , sin "\alpha" = cos "\beta" , or sin "\beta =" cos "\alpha"


1) Let us look at the last 2 cases (sin "\alpha" = cos "\beta" , or sin "\beta" = cos "\alpha" )

sin "\alpha" = cos "\beta" "\rArr \alpha=90^o-\beta \rArr" sin "\beta" = cos "\alpha"

So this cases are actually the same

We know that "\alpha=4x+10" and "\beta = 10x + 10"

We can substitute these into "\alpha=90^o-\beta"

"\alpha=90^o-\beta \rArr 4x+10=90-(10x+10) \rArr 4x+10=90-10x-10"

"(4x+10)+10x-10=(90-10x-10)+10x-10"

"14x=70 \rArr x=5 \rArr \alpha = 4*5+10=30^o \rArr \beta=90^o-\alpha=90^o-30^o=60^o"



2) Let us look at the case sin "\alpha" = cos "\alpha" .

This means that "\alpha=45^o \rArr 4x+10=45^o \rArr 4x=35^o \rArr 10x=35*\frac{10}{4}=87^o.5"

"\beta=10x+10=97^o.5"


3) Let us look at the case sin "\beta =" cos "\beta"

This means that "\beta = 45^o \rArr 10x+10=45^o \rArr 10x=35^o \rArr 4x=35^o\frac{4}{10}=14^o"

"\alpha=4x+10=24^o"